Subscribing to the theory that life is indeed a beach, the residents of La Playa spend all of their money on three things: Every year, they collectively buy 225 bathing suits, 620 tubes of sunscreen, and 380 beach towels Item (amount purchased) Price 2013 () Price 2014 () Price 2015 () Price 2016 ($) Bathing suits (225 Sunscreen (620 Beach towels (380 $9.00 4.50 4.00 $10.00 4.00 5.25 $14.00 4.00 6.00 $16.00 5.00 8.00 Use the data in the table above to calculate the following. a. The total cost of this basket each year from 2013 through 2016 Instructions: Round your answers to the nearest whole number Price 2013 () Price 2014 ( Price 2015 () Price 2016 () Total cost of basket b. How much the cost of this basket has changed from year to year in percentage terms Instructions: Round your answers to one decimal place 2013-2014 2014-2015 2015-2016 Percentage change in total cost of basket
Answer #1

(a)

Total cost of basket, 2013 ($) = 225 x 9 + 620 x 4.5 + 380 x 4 = 2025 + 2790 + 1520 = 6335

Total cost of basket, 2013 ($) = 225 x 10 + 620 x 4 + 380 x 5.25 = 2250 + 2480 + 1995 = 6725

Total cost of basket, 2013 ($) = 225 x 14 + 620 x 4 + 380 x 6 = 3150 + 2480 + 2280 = 7910

Total cost of basket, 2013 ($) = 225 x 16 + 620 x 5 + 380 x 8 = 3600 + 3100 + 3040 = 9740

(b)

% Change, 2013-14 = (6725 / 6335) - 1 = 1.062 - 1 = 0.062 = 6.2%

% Change, 2014-15 = (7910 / 6725) - 1 = 1.176 - 1 = 0.176 = 17.6%

% Change, 2015-16 = (9740 / 7910) - 1 = 1.231 - 1 = 0.231 = 23.1%